### v1

@@ Abstract
Smart assistive technologies such as sensor-based footwear and walking aids offer promising opportunities to support rehabilitation through real-time feedback and patient-centered monitoring. However, most orthotic devices remain passive and lack integrated sensing or feedback functionalities, while existing research often focuses on isolated prototypes rather than cohesive, interactive systems. In this work, we present the design and implementation of a novel modular sensor system that combines a smart foot orthosis with an instrumented forearm crutch. The system integrates plantar pressure and motion sensing, vibrotactile feedback, and wireless communication via a smartphone application. We conducted an experimental user study with eight participants to validate the feasibility of the smart foot orthosis for mobile gait detection, explore the potential of haptic feedback for user interaction, and assess the usability of the accompanying mobile health application. Our work contributes to the field of smart assistive technology in rehabilitation and prevention by demonstrating a functional and comprehensive system. We further discuss system limitations, outline potential application scenarios, and provide recommendations for future development and clinical integration.
Index Terms:
Assistive Technology, Foot Health, Gait Analysis, Haptic Feedback, Instrumented Walking Aid, mHealth Application, Rehabilitation, Smart Insole, Smart Orthosis, Wearable Sensors.
@@ I Introduction
Gait rehabilitation plays an important role in the recovery process after acute foot injuries and in the management of chronic foot-related conditions. A variety of assistive devices are used to support these patients in their daily mobility, including orthopedic devices such as foot orthoses, insoles, bandages, and pressure-relieving footwear [1]. In addition, mobility aids such as crutches, walking sticks, or canes are commonly used for temporary or long-term walking impairments [2]. While these devices are essential in stabilizing gait and reducing mechanical stress [3, 4], they are typically standardized and require customization to achieve optimal recovery outcomes [5]. These limitations underscore the need for more individualized and adaptive orthopedic solutions that combine health monitoring with interactive feedback.
@@
In recent years, the integration of wearable electronics has contributed to various developments such as smart insoles, socks, and shoes equipped with sensor technologies for gait analysis and health monitoring [6, 7, 8]. In contrast to conventional stationary systems, these mobile solutions enable gait analysis outside the clinical settings and open up new possibilities for continuous monitoring of spatial, temporal, and spatio-temporal parameters [9]. These systems frequently rely on force sensing resistors and inertial measurement units to assess parameters such as plantar pressure distribution and gait patterns (e.g., step cadence and foot strike patterns) [10]. Such technologies show promise in areas such as fall risk assessment, biofeedback interventions, and the prevention of foot ulcers [11]. In particular, monitoring plantar pressure is crucial for patients with diabetic foot syndrome (DFS), where elevated sole pressure can impair wound healing and increase the risk of ulcer formation [12]. They also hold potential for injury prevention and early pathology detection [13].
Despite this progress, smart gait monitoring systems are still largely limited to insole-based or shoe-integrated solutions, which are mostly developed for athletic performance metrics or general health monitoring [13]. Their application in gait rehabilitation, especially for individuals with mobility impairments, remains limited. In particular, orthopedic devices such as ankle-foot orthoses, which are used to stabilize gait and support recovery, still lack the integration of embedded sensor and feedback technologies. Further development is necessary to fully realize their potential in rehabilitation contexts. For individuals with foot conditions, integrating feedback modalities (such as haptic alerts triggered by pressure overload) could enhance the benefits of smart foot orthosis (SFO) development [14]. For example, diabetic neuropathy in the lower extremities impairs the sensation of pain and limits the detection of excessive plantar pressure [15]. As a result, injuries can remain unnoticed, which increases the risk of serious deformities such as Charcot foot [16]. Providing patients with internal feedback could positively impact gait recovery. Despite this potential, only a few conceptual approaches have explored the integration of embedded electronics into AFOs [17, 18, 19] for measuring physiological parameters. Existing research is mostly limited to isolated prototype solutions. Moreover, no comprehensive system has yet been presented that integrates pressure and gait data acquisition with haptic feedback and real-time communication between a smart foot orthosis, a corresponding walking aid, and a mobile application.
@@
To address the identified research gap, we present a novel multimodal system for interactive gait monitoring. The system comprises three key components: (1) a SFO equipped with IMU and FSR sensors as well as a vibration feedback module, (2) a sensor-integrated walking aid that communicates with the orthosis and includes its own IMU and haptic actuators, and (3) a mobile smartphone application for health monitoring and feedback interaction. The system architecture enables real-time communication across all components and supports personalized health monitoring.
Based on this concept, the following research questions are addressed:
RQ1: How accurate are the plantar pressure and gait measurements of the developed SFO compared to a clinical reference system?
@@
RQ2: How do different haptic feedback types (continuous vs. pulsed), implemented in a smart foot orthosis and a smart walking aid, affect user perception regarding noticeability, intrusiveness, usefulness, and intended use?
RQ3: How usable is the accompanying mobile application for visualizing gait and pressure data from a user experience perspective?
To assess the technical validity of the sensor data, we compare gait parameters recorded by the SFO with a clinical reference system using a Zebris treadmill for gait analysis. This comparison allows us to evaluate the gait data reliability of the embedded sensor system and explore the potential for future use. By demonstrating the feasibility of our system, we aim to advance the field of intelligent assistive technologies in gait rehabilitation and prevention. This includes the validation of sensor-based measurements from the SFO, the evaluation of user perception regarding haptic feedback types in orthopedic devices, and the usability assessment of the accompanying mobile application for health monitoring and interaction. Our findings serve as a basis for future patient-centered evaluations focusing on the effectiveness of multimodal feedback. Building on the development and validation of our integrated system, this work contributes the following key innovations to the field of sensor-driven and interactive rehabilitation technologies:
  • Design and implementation of a SFO combining plantar pressure and gait sensing with integrated vibrotactile feedback.
  • Technical validation of the SFOs gait analysis functionality through comparison with a clinical reference system to assess measurement reliability.
  • Development and user experience evaluation of a sensor-based walking aid enabling real-time haptic feedback via communication with the SFO and a mobile app.
  • Implementation and usability testing of a mobile application for real-time health monitoring, configuration of feedback modalities, and cross-device data visualization.
II Related Work
@@
This section reviews related research in the areas of smart footwear for mobile gait analysis and health monitoring (subsection II-A), foot augmentation with multimodal feedback for patient–device interaction (subsection II-B), and sensor-integrated walking aids for mobility support (subsection II-C). We conclude by summarizing current research gaps and outlining our contributions to the field.
@@
II-A Smart Footwear for Mobile Gait Analysis and Monitoring Health Parameters
@@
The research field of sensor technology demonstrates that different types of sensors are used for measuring continuous physiological data depending on the application [13]. Based on the device type, sensor technologies can be integrated into various systems such as socks/textiles [20, 21, 22, 23, 24], shoes [25, 26, 27, 28, 29], insoles [30, 7, 31, 8, 32], or orthoses [17, 18, 33, 34, 35]. In particular, smart insoles are developed for a broad range of applications, ranging from ambulatory gait assessment and home-based monitoring [36, 37], to gait rehabilitation after stroke [38], fall risk detection [39], and activity classification using machine learning techniques [40, 41].
@@
A systematic literature review by Subramaniam et al. [13] identified two primary health-related functions of smart insoles: (1) the measurement of plantar pressure distribution and (2) the analysis of gait and activity patterns. Plantar pressure data are typically acquired using integrated FSR, for the assessment of foot stability, balance, body weight distribution, and early detection of pressure overload conditions. A review by Razak et al. [42] recommended the use of 15 FSR sensors to ensure sufficient coverage of plantar pressure distribution according to the individual foot anatomy. This recommendation is based on an anatomical segmentation of the foot sole into 15 regions: heel (areas 1–3), midfoot (4–5), metatarsals (6–10), and toes (11–15), which support most of the body load [43]. However, the individual shoe size and foot anatomy needs to be considered for proper sensor placement to prevent measurement errors [44].
@@
Furthermore, the use of IMU sensors (a combination of accelerometers and gyroscopes) enables the measurement of gait characteristics such as step length, cadence, walking speed, and swing phase duration [45]. These systems are used for performance tracking of athletes [13] or to enable mobile gait analysis [46]. A single IMU sensor embedded in the insole can provide sufficient data for this purpose [47]. However, the literature reports inconsistencies regarding the optimal sensor placement for precise data extraction, as signal drift and measurement errors may occur depending on the location of the sensor [48]. As a result, calibration remains a critical challenge in smart insole development [13]. To minimize mechanical load and improve signal quality, placement in the midfoot region is generally recommended [13].
@@
Beyond IMU and FSR-based gait analysis, additional sensing modalities have been explored to extend the functionality of wearable systems. These include Global Positioning System (GPS) tracking for spatial gait analysis [49], foot temperature monitoring for early detection of DFS [11, 50], and physiological measurements such as heart rate via photoplethysmography (PPG) sensors [51] and muscle activity using electromyography (EMG) electrodes in textile-based wearables [21]. However, PPG and EMG-based approaches remain largely limited to research prototypes and have not yet been adopted in commercial devices. Other studies have also investigated sensing technologies for obstacle detection, including ultrasonic sensors, ToF cameras, and radar-based systems, particularly in applications for visually impaired users [52]. While these alternative sensing modalities offer promising extensions for specific use cases, the majority of current smart insole systems rely on the integration of IMU and FSR sensors. Due to its efficiency and widespread adoption, the combination is frequently used in smart insole systems [53]. This configuration enables accurate gait detection [54, 55] and simultaneous acquisition of kinetic (pressure) and kinematic (motion) data [10, 56]. Typically, data from IMU and FSR sensors are acquired using a portable microcontroller and transmitted wirelessly (e.g., via Bluetooth) to an external graphical user interface (GUI[36, 57].
@@
To ensure the reliability of gait-related measurements derived from wearable systems, their performance must be validated against established reference standards. A review by Hutabarat et al. [9] emphasized that such wearable gait systems need to be validated against proven standard gait measurement systems to ensure data accuracy. For this purpose, instrumented treadmills and motion capture systems represent the gold standard for validating gait analysis systems [58, 9]. This approach has already been applied in several validation studies, for instance by comparing wearable systems to a stationary Zebris treadmill [59] or GAITRite force platforms [60, 61]. Overall, recent research highlights the importance of comparative validation studies with established gait analysis systems to strengthen the credibility and reliability of findings in wearable gait assessment [62]. In conclusion, despite considerable advances in this research field, key challenges remain regarding data accuracy, system validation, and participant-centered evaluation.
@@
II-B Foot Augmentation with Multimodal Feedback for Patient Device Interaction
@@
Smart wearable systems such as insoles, socks, or orthoses are often connected to external devices (e.g., smartphone applications, smartwatches, or other interfaces) to enable continuous mobile health monitoring and interactive feedback [13]. A systematic review by Saboor et al. [63] highlighted the growing potential of smartphone applications for mobile gait analysis, particularly in remote health contexts, while also noting ongoing challenges related to accessible and user-friendly user interface (UI) design. In this context, Resch et al. [64] developed a mobile Health (mHealth) application for smart insoles on the basis of focus groups with patients and experts. In a follow-up study, functional mockups were evaluated in a usability test with 30 participants, which confirmed the user-friendliness and provided recommendations for future design improvements [65]. The resulting design recommendations, integrated features, and overall app concept provide a valuable basis for the continued development of smart insole applications and the integration of feedback modalities.
@@
Beyond interface design, wearable systems increasingly incorporate real-time feedback mechanisms (through visual, acoustic, or haptic signals) typically via smartphone apps or connected devices to support user interaction. For instance, vibration cues have been applied to indicate pressure overload [66]. It was shown that vibration feedback applied within the shoe can significantly influence gait behavior in healthy participants [67], and vibrotactile biofeedback triggered by FSR input reduced overpronation duration by 50% [68]. Furthermore, auditory feedback has also been investigated, such as acoustic signals triggered by in-sock pressure sensors for gait correction [69], or auditory stimulation via insoles for patients with Parkinson’s disease [70]. These approaches are particularly relevant for patients with impaired peripheral sensation, such as individuals with DFS, for whom smartwatch-based alerts and real-time gait correction via smart insoles have been proposed [71]. Several studies have applied multimodal feedback approaches and shown that they influence gait symmetry, posture and pressure distribution [72, 67, 68, 73]. In an experimental study, Redd et al. [72] compared visual, acoustic, and vibrotactile feedback modalities in healthy participants and found that both visual and haptic cues significantly improved gait symmetry. Despite the demonstrated potential of smartphone applications and feedback modalities, further validation is needed regarding user perception and system effectiveness.
II-C Smart Walking Aids and Applied Sensor Technologies

### v2

@@ Abstract
Smart assistive technologies such as sensor-based footwear and walking aids offer promising opportunities for gait rehabilitation through real-time feedback and patient-centered monitoring. While biofeedback applications show great potential, current research rarely explores integrated closed-loop systems with device- and modality-specific feedback. In this work, we present a modular sensor-based system combining a smart foot orthosis and an instrumented forearm crutch to deliver real-time vibrotactile biofeedback. The system integrates plantar pressure and motion sensing, vibrotactile feedback, and wireless communication via a smartphone application. We conducted a user study with eight participants to validate the system’s feasibility for mobile gait detection and app usability, and to evaluate different vibrotactile feedback types across the orthosis and forearm crutch. The results indicate that pattern-based vibrotactile feedback was rated as more useful and suitable for regular use than simple vibration alerts. Moreover, participants reported clear perceptual differences between feedback delivered via the orthosis and the forearm crutch, indicating device-dependent feedback perception. The findings highlight the relevance of feedback strategy design beyond hardware implementation and inform the development of user-centered haptic biofeedback systems.
keywords:
assistive technology; foot health; gait analysis; haptic feedback; instrumented walking aid; mHealth application; rehabilitation; smart insole; smart orthosis; wearable sensors
@@ 1 Introduction
Gait rehabilitation plays an important role in the recovery process after acute foot injuries and in the management of chronic foot-related conditions. A variety of assistive devices are used to support these patients in their daily mobility, including orthopedic devices such as foot orthoses, insoles, bandages, and pressure-relieving footwear Barbara and Grobelna (2022). In addition, mobility aids such as crutches, walking sticks, or canes are commonly used for temporary or long-term walking impairments Bertrand et al. (2017). While these devices are essential in stabilizing gait and reducing mechanical stress Burns et al. (2009); de Jong et al. (2023), they are typically standardized and require customization to achieve optimal recovery outcomes Arch and Stanhope (2017), as they can measurably affect gait Resch and Carroll (2025). These limitations underscore the need for more individualized and adaptive orthopedic solutions that combine health monitoring with interactive feedback.
@@
In recent years, the integration of wearable electronics has contributed to various developments such as smart insoles, socks, and shoes equipped with sensor technologies for gait analysis and health monitoring Salis et al. (2023); Almuteb et al. (2022); Okoduwa et al. (2024). In contrast to conventional stationary systems, these mobile solutions enable gait analysis outside clinical settings and open up new possibilities for continuous monitoring of spatial, temporal, and spatio-temporal parameters Hutabarat et al. (2021). These systems frequently rely on force sensing resistors and inertial measurement units to assess parameters such as plantar pressure distribution and gait characteristics (e.g., step cadence and foot strike patterns) Kim et al. (2018). Such technologies show promise in areas such as fall risk assessment, biofeedback interventions, and the prevention of foot ulcers Grech et al. (2022). In particular, monitoring plantar pressure is crucial for patients with diabetic foot syndrome (DFS), where elevated sole pressure can impair wound healing and increase the risk of ulcer formation Frykberg et al. (1998). They also hold potential for injury prevention and early pathology detection Subramaniam et al. (2022).
However, orthopedic devices such as ankle-foot orthoses, which are used to stabilize gait and support recovery, still rarely combine embedded sensing and feedback within an interactive system. Integrating feedback modalities (such as haptic alerts triggered by pressure overload) could enhance the benefits of smart foot orthosis (SFO) development, particularly for individuals with foot conditions Resch et al. (2024). For example, diabetic neuropathy in the lower extremities impairs the sensation of pain and limits the detection of excessive plantar pressure Alam et al. (2017), which can lead to unnoticed injuries and increase the risk of serious deformities such as Charcot foot Rogers et al. (2011). Providing patients with intrinsic feedback could therefore positively impact gait recovery. While several conceptual prototypes have integrated electronics into orthoses Finco et al. (2023); Park et al. (2023); Zhou et al. (2021), most solutions remain technically focused, without addressing how haptic feedback should be designed or delivered. There is a lack of empirical evidence on how feedback modality in smart footwear and walking aids influences user perception, perceived effectiveness, and intended use. This is critical, as unclear or intrusive feedback can reduce acceptance even when technically accurate.
@@
To address the identified research gap, we developed a modular sensor system to investigate the influence of haptic feedback modality and device-dependent feedback on user perception in smart footwear and walking aids. The system comprises three key components: (1) an SFO equipped with IMU and FSR sensors as well as a vibration feedback module, (2) a sensor-integrated walking aid that communicates with the orthosis and includes its own IMU and haptic actuators, and (3) a mobile smartphone application for health monitoring and feedback interaction. The system architecture enables real-time communication across all components and provides the basis for evaluating closed-loop vibrotactile biofeedback strategies. To ensure a valid basis for user perception testing, we examined the system’s feasibility for gait detection and the usability of the mobile application from a user interaction perspective. These validation objectives guided the following supporting research questions:
  • SRQ1 (System Feasibility): How accurately does the developed SFO measure plantar pressure and gait parameters compared to a clinical reference system, as a foundation for reliable real-time multimodal feedback?
  • SRQ2 (Application Usability): How do users evaluate the usability of the accompanying mobile application for interpreting gait and pressure information within the integrated system?
Based on this, the core objective of this work is to investigate feedback strategy design for user-centered biofeedback, addressing the main research question (MRQ):
@@
  • MRQ (Feedback Evaluation): How do different vibrotactile feedback types (continuous vs. pattern-based), delivered through the orthosis and the walking aid, influence user perception in terms of noticeability, intrusiveness, perceived usefulness, and intended use?
This work contributes to the field of assistive technology by introducing a novel closed-loop, sensor-driven haptic feedback mechanism in a sensorized foot orthosis and instrumented forearm crutch, enabling real-time feedback that responds to gait dynamics and pressure distribution. Our key research findings provide empirical insights that users significantly prefer pattern-based over simple vibrotactile feedback notifications and that device location influences user perception. These results offer concrete recommendations for designing effective, user-centered haptic rehabilitation systems and can serve as a basis for future patient-centered evaluations of multimodal feedback in human-computer interaction (HCI).
2 Related Work
@@
This section reviews related research in the areas of smart footwear for mobile gait analysis and health monitoring (Section 2.1), foot augmentation with multimodal feedback for patient–device interaction (Section 2.2), and sensor-integrated walking aids for mobility support (Section 2.3). We conclude by summarizing current research gaps and outlining our contributions to the field.
@@
2.1 Smart Footwear for Mobile Gait Analysis and Monitoring Health Parameters
@@
The research field of sensor technology demonstrates that different types of sensors are used for measuring continuous physiological data depending on the application Subramaniam et al. (2022). Based on the device type, sensor technologies can be integrated into various systems such as socks/textiles Najafi et al. (2017); Isezaki et al. (2019); Zhang et al. (2020); Drăgulinescu et al. (2020); Semjonova et al. (2022), shoes Morley et al. (2001); Eskofier et al. (2017); Sunarya et al. (2020); Gokalgandhi et al. (2020); Ullrich et al. (2020), insoles Lin et al. (2016); Almuteb et al. (2022); Xu et al. (2012); Okoduwa et al. (2024); Duong et al. (2020), or orthoses Finco et al. (2023); Park et al. (2023); Shefa et al. (2023, 2024); Han et al. (2024). In particular, smart insoles are developed for a broad range of applications, ranging from ambulatory gait assessment and home-based monitoring Mustufa et al. (2015); Roth et al. (2018), to gait rehabilitation after stroke Farid et al. (2021), fall risk detection Chen et al. (2021), and activity classification using machine learning techniques D’arco et al. (2024); Djamaa et al. (2020).
@@
A systematic literature review by Subramaniam et al. Subramaniam et al. (2022) identified two primary health-related functions of smart insoles: (1) the measurement of plantar pressure distribution and (2) the analysis of gait and activity patterns. Plantar pressure data are typically acquired using integrated FSRs for the assessment of foot stability, balance, body weight distribution, and early detection of pressure overload conditions. A review by Razak et al. Razak et al. (2012) recommended the use of 15 FSR sensors to ensure sufficient coverage of plantar pressure distribution according to the individual foot anatomy. This recommendation is based on an anatomical segmentation of the foot sole into 15 regions: heel (areas 1–3), midfoot (4–5), metatarsals (6–10), and toes (11–15), which support most of the body load Shu et al. (2010). However, the individual shoe size and foot anatomy need to be considered for proper sensor placement to prevent measurement errors Senanayake and Senanayake (2010).
@@
Furthermore, the use of IMU sensors (a combination of accelerometers and gyroscopes) enables the measurement of gait characteristics such as step length, cadence, walking speed, and swing phase duration Gujarathi and Bhole (2019). These systems are used for performance tracking of athletes Subramaniam et al. (2022) or to enable mobile gait analysis Aggarwal et al. (2018). A single IMU sensor embedded in the insole can provide sufficient data for this purpose Sabatini et al. (2005). However, the literature reports inconsistencies regarding the optimal sensor placement for precise data extraction, as signal drift and measurement errors may occur depending on the location of the sensor Küderle et al. (2022). As a result, calibration remains a critical challenge in smart insole development Subramaniam et al. (2022). To minimize mechanical load and improve signal quality, placement in the midfoot region is generally recommended Subramaniam et al. (2022).
@@
Beyond IMU and FSR-based gait analysis, additional sensing modalities have been explored to extend the functionality of wearable systems. These include Global Positioning System (GPS) tracking for spatial gait analysis Zhang and Poslad (2014), foot temperature monitoring for early detection of DFS Grech et al. (2022); Matijevich et al. (2024), and physiological measurements such as heart rate via photoplethysmography (PPG) sensors Hong and Park (2018) and muscle activity using electromyography (EMG) electrodes in textile-based wearables Isezaki et al. (2019). However, PPG and EMG-based approaches remain largely limited to research prototypes and have not yet been adopted in commercial devices. Other studies have also investigated sensing technologies for obstacle detection, including ultrasonic sensors, ToF cameras, and radar-based systems, particularly in applications for visually impaired users Joseph et al. (2023). While these alternative sensing modalities offer promising extensions for specific use cases, the majority of current smart insole systems rely on the integration of IMU and FSR sensors. Due to its efficiency and widespread adoption, the combination is frequently used in smart insole systems Manna et al. (2023). This configuration enables accurate gait detection Prasanth et al. (2021); Han et al. (2019) and simultaneous acquisition of kinetic (pressure) and kinematic (motion) data Kim et al. (2018); Manupibul et al. (2023). Typically, data from IMU and FSR sensors are acquired using a portable microcontroller and transmitted wirelessly (e.g., via Bluetooth) to an external graphical user interface (GUIMustufa et al. (2015); Benocci et al. (2009).
@@
To ensure the reliability of gait-related measurements derived from wearable systems, their performance must be validated against established reference standards. A review by Hutabarat et al. Hutabarat et al. (2021) emphasized that such wearable gait systems need to be validated against proven standard gait measurement systems to ensure data accuracy. For this purpose, instrumented treadmills and motion capture systems represent the gold standard for validating gait analysis systems Jakob et al. (2021); Hutabarat et al. (2021). This approach has already been applied in several validation studies, for instance by comparing wearable systems to a stationary Zebris treadmill Donath et al. (2016) or GAITRite force platforms Jagos et al. (2017); Raghuvanshi et al. (2024). Overall, recent research highlights the importance of comparative validation studies with established gait analysis systems to strengthen the credibility and reliability of findings in wearable gait assessment Santos et al. (2024). In conclusion, despite considerable advances in this research field, key challenges remain regarding data accuracy, system validation, and participant-centered evaluation.
@@
2.2 Foot Augmentation with Multimodal Feedback for Patient Device Interaction
@@
Smart wearable systems such as insoles, socks, or orthoses are often connected to external devices (e.g., smartphone applications, smartwatches, or other interfaces) to enable continuous mobile health monitoring and interactive feedback Subramaniam et al. (2022). A systematic review by Saboor et al. Saboor et al. (2020) highlighted the growing potential of smartphone applications for mobile gait analysis, particularly in remote health contexts, while also noting ongoing challenges related to accessible and user-friendly user interface (UI) design. In this context, Resch et al. Resch et al. (2023) developed a mobile Health (mHealth) application for smart insoles on the basis of focus groups with patients and experts. In a follow-up study, functional mockups were evaluated in a usability test with 30 participants, which confirmed the user-friendliness and provided recommendations for future design improvements Resch et al. (2024). The resulting design recommendations, integrated features, and overall app concept provide a valuable basis for the continued development of smart insole applications and the integration of feedback modalities.
@@
Beyond interface design, wearable systems increasingly incorporate real-time feedback mechanisms (through visual, acoustic, or haptic signals) typically via smartphone apps or connected devices to support user interaction. For instance, vibration cues have been applied to indicate pressure overload Crea et al. (2015). It was shown that vibration feedback applied within the shoe can significantly influence gait behavior in healthy participants Oates et al. (2017), and vibrotactile biofeedback triggered by FSR input reduced overpronation duration by 50% Berengueres et al. (2014). Furthermore, auditory feedback has also been investigated, such as acoustic signals triggered by in-sock pressure sensors for gait correction Biesmans and Markopoulos (2020), or auditory stimulation via insoles for patients with Parkinson’s disease Pacilli et al. (2016). These approaches are particularly relevant for patients with impaired peripheral sensation, such as individuals with DFS, for whom smartwatch-based alerts and real-time gait correction via smart insoles have been proposed Ma et al. (2016). Several studies have applied multimodal feedback approaches and shown that they influence gait symmetry, posture and pressure distribution Redd and Bamberg (2012); Oates et al. (2017); Berengueres et al. (2014); Elvitigala et al. (2019). In an experimental study, Redd et al. Redd and Bamberg (2012) compared visual, acoustic, and vibrotactile feedback modalities in healthy participants and found that both visual and haptic cues significantly improved gait symmetry. Despite the demonstrated potential of smartphone applications and feedback modalities, further validation is needed regarding user perception and system effectiveness.
2.3 Smart Walking Aids and Applied Sensor Technologies